OpenStudy (anonymous):

Solve for y: 2/3y + 4 = 6

OpenStudy (tkhunny):

Have you considered subtracting 4?

OpenStudy (anonymous):

yes, and I got the equation: \[\frac{ 2 }{ 3 }y\]=2

OpenStudy (tkhunny):

Okay, now what? How do we get that 2/3 off of there?

OpenStudy (anonymous):

divide or multiply by 2/3?

OpenStudy (tkhunny):

Divide be 2/3 or multiply by 3/2. You must pick. The choice is yours. Generally, multiplication is easier.

OpenStudy (anonymous):

so the equation left would be y=3/2+2

OpenStudy (tkhunny):

No, how do we get addition on the right hand side? \(\dfrac{2}{3}y = 2\) \(\dfrac{3}{2}\cdot\dfrac{2}{3}y = \dfrac{3}{2}\cdot2\) \(y = \dfrac{3}{2}\cdot 2\) Simplify the right hand side.

OpenStudy (anonymous):

@tkhunny

OpenStudy (tkhunny):

Did you simplify it?

OpenStudy (shaik0124):